OpenStudy (anonymous):

The integral f(x) = 3x/5 -2 from 0 to 5 should give the answer 3. I get -2,5 My attempt: F(x) = 3x² / 10 - 2x F(x) = 3*5² / 10 - (2 * 5) = -2,5

OpenStudy (anonymous):

As you can see, along our interval, e^(x) > e^(-3x), so our integral will be: (e^(x) - e^(-3x)) * dx, from x = 0 to x = ln(5) The integral of e^(x) * dx is e^(x) The integral of e^(-3x) * dx is (-1/3) * e^(-3x) e^(x) - (-1/3) * e^(-3x) + C => e^(x) + (1/3) * e^(-3x) + C From 0 to ln(5) e^(ln(5)) - e^(0) + (1/3) * e^(-3 * ln(5)) - (1/3) * e^(-3 * 0) => 5 - 1 + (1/3) * 5^(-3) - (1/3) => 4 - 1/3 + (1/3) * (1/125) => 11/3 + 1/375 => (11 * 125 + 1) / 375 => (1375 + 1) / 375 => 1376/375 => 3.6693333333333333333333333333333

OpenStudy (anonymous):

is that goood

OpenStudy (amistre64):

integrate 3x/5 -2 ; [0,5] since this is a poly, the 0 is pointless to us; this IS a poly right? 3(5)^2/10 - 2(5) 3(25)/10 - 10 75/10 - 10 = 7.5 - 10 = -2.5

OpenStudy (amistre64):

0 to 5, or 0 to ln(5)?

OpenStudy (anonymous):

hmm, I must be missing something; the answer should be 3... :/

OpenStudy (anonymous):

0 to 5

OpenStudy (anonymous):

I don't know what you mean by a poly

OpenStudy (amistre64):

of the form\[c_0+c_1x+c_2x^2+...+c_nx^n\] as opposed to a rational expression of some sort

OpenStudy (amistre64):

poly is short for polynomial

OpenStudy (anonymous):

hmm, no this is a straight line, so y = kx + m

OpenStudy (amistre64):

a poly of degree 1 is a straight line

OpenStudy (amistre64):

can you scan the problem? maybe your misreading something?

OpenStudy (amistre64):

....or you might be looking at the wrong answer key, or the answer key may even have a wrong solution. That happens from time to time

OpenStudy (anonymous):

I'm thinking the answer key seems to be wrong. There isn't much to misread and you seem to reach the same answer as me. So, I'm going to assume the answer 3 is wrong. Help appreciated. :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+3x%2F5-2+dx%2C+from+0..5 just as a dbl chk :) good luck

OpenStudy (anonymous):

thx